# Minimal Subsampled Rank-1 Lattices for Multivariate Approximation with Optimal Convergence Rate

### v2

@@ Abstract
MSC2020: 41A25, 94A20
1 Introduction
This paper mainly deals with the question of approximating functions from partial rank-1 lattices. Rank-1 lattices have their origin as a Quasi-Monte Carlo (QMC) method and have been thoroughly studied in various settings ever since, see e.g., [Sloan94, kaemmererdiss, KaPoVo13, PlPoStTa18, CKNS20, DKP22]. The lattice point sets are given by $\bm{X}=\{\tfrac{k}{n}\bm{z}\bmod{}1\}_{k\in\{0,\dots,n-1\}}$, where $\bm{z}\in\mathds{Z}^{d}$ is known as the generating vector and it controls the quality of the point set, and mod $\bmod\,1$ denotes a component-wise fractional part of a vector. The study of partial or subsampled rank-1 lattices has only begun recently, cf. [BKPU22]. Such point sets have the form
@@ 1 Introduction
An example is depicted in Figure 1.
$\bm{X}$$\bm{X}_{J}$efer to caption
Figure 1: Left: rank-1 lattice with generating vector $\bm{z}=(1,21)$ and lattice size $n=55$.
Right: subsampled rank-1 lattice with $|J|=\lceil\sqrt{n}\,\log\sqrt{n}\,\rceil=15$.
The structure of lattice point sets allows for the application of fast Fourier methods when approximating functions whilst the subsampling improves the approximation properties, cf. [BSU23]. We investigate the worst-case setting, where the point set and algorithm are fixed and are used to approximate every function from a given function class $H$. The worst-case error for a general algorithm $A$ measured in $L_{2}$-norm is defined by
@@
    $e^{\mathrm{wor-app}}(A_{\mathcal{A}}^{\bm{X}})\lesssim n^{-\alpha/2+\varepsilon}\,,$ (1.1)
    where $\varepsilon>0$, $A_{\mathcal{A}}^{\bm{X}}$ is the “classical lattice algorithm” (see (2.1) below), and $A\lesssim B$ means there exists a constant $C>0$ independent of $n$ such that $A\leq CB$. When $\alpha$ is an integer it is the number of available square-integrable derivatives in each coordinate.
@@
In this paper we connect both approaches by finding a frequency index set for which a lattice is reconstructing in Theorem 3.2 below. These frequencies solely depend on the worst-case error achievable using any algorithm and the lattice in question. In that way, lattices constructed according the first approach can be investigated with respect to their reconstructing property, and tools from the second approach become available.
@@
$e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})\lesssim|J|^{-\alpha}(\log|J|)^{(d-1)\alpha+\frac{1}{2}}\quad\text{with}\quad n\sim|J|^{\frac{2}{1-1/(2\alpha)}}\,,$ (1.2)
where $S_{\mathcal{B}}^{\bm{X}_{J}}$ is the “least squares approximation” to be defined precisely in (4.1) below and $A\sim B$ means $A\lesssim B$ and $B\lesssim A$ holds simultaneously.
This method achieves the optimal polynomial order $|J|^{-\alpha+\varepsilon}$ up to $\varepsilon>0$ in terms of sampling complexity. However, the computational complexity is determined by $n$. Since an algorithm using a subsampled lattice is inherently still using the lattice of size $n$, the lower bound from [BKUV17] still applies, and achieving the rate $|J|^{-\alpha+\varepsilon}$ implies the lower bound
@@
With respect to the size of $|J|$, (1.2) is off by a polynomial order depending on the smoothness $\alpha$.
In this paper, in Theorem 5.3 below, we show that we can go arbitrarily close to the squared number of points independently of the smoothness, i.e., we have
$e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})\lesssim\Big{(}\frac{|J|}{\log|J|}\Big{)}^{-\alpha+\varepsilon}\quad\text{with}\quad\Big{(}\frac{|J|}{\log|J|}\Big{)}^{2{\sqrt{1-\varepsilon/\alpha}}}\lesssim\,n\lesssim\Big{(}\frac{|J|}{\log|J|}\Big{)}^{\frac{2}{\sqrt{1-\varepsilon/\alpha}}}\,,$
@@
This paper is organized as follows. We start by introducing the classical lattice algorithm and the kernel method in Section 2. Then we investigate the projection property of approximation algorithms in Section 3, where we also prove a connection between the worst-case error and the reconstructing property of lattices in Theorem 3.2, which appears to be a new result. In Section 4 we show general error bounds for the randomly subsampled lattice, which we apply to the Korobov space setting in Section 5. After some comments on different implementations in Section 6, we present some numerical experiments in Section 7 and give concluding remarks in Section 8.
@@
[…]
2 Classical lattice algorithm versus kernel method
@@
$A_{\mathcal{A}}^{\bm{X}}f\coloneqq\sum_{\bm{h}\in\mathcal{A}}\Big{(}\underbrace{\frac{1}{n}\sum_{k=0}^{n-1}f(\tfrac{k}{n}\bm{z}\bmod{}1)\exp(-2\pi\mathrm{i}\tfrac{k}{n}\langle\bm{h},\bm{z}\rangle)}_{\eqqcolon\,\widehat{g}_{\bm{h}}}\Big{)}\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\,,$ (2.1)
see e.g., [KSW06, KaPoVo13, kaemmererdiss, PlPoStTa18, CKNS20, DKP22]. To get a feeling for this method, we have a look at the Fourier coefficients of the approximation $A_{\mathcal{A}}^{\bm{X}}f\eqqcolon g=\sum_{\bm{h}\in\mathcal{A}}\widehat{g}_{\bm{h}}\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)$ in terms of the Fourier coefficients of $f$. By the character property $\frac{1}{n}\sum_{k=0}^{n-1}\exp(2\pi\mathrm{i}m\tfrac{k}{n})=\delta_{m\equiv_{n}0}$, we have for $\bm{h}\in\mathcal{A}$
@@
$=\sum_{\begin{subarray}{c}\bm{\ell}\in\mathds{Z}^{d}\\ \langle\bm{\ell},\bm{z}\rangle\equiv_{n}\langle\bm{h},\bm{z}\rangle\end{subarray}}\!\!\!\!\!\!\widehat{f}_{\bm{\ell}}\,.$ (2.2)
@@
$f\in H=\Big{\{}f\colon\mathds{T}^{d}\to\mathds{C}:\|f\|_{H}^{2}=\sum_{\bm{h}\in\mathds{Z}^{d}}r(\bm{h})\,|\widehat{f}_{\bm{h}}|^{2}<\infty\Big{\}}\,,$ (2.3)
where $r(\bm{h})\to\infty$ for $\|\bm{h}\|_{2}\to\infty$. Specific choices of the function $r(\cdot)$ lead to e.g., the Korobov spaces in Section 5. Obviously there are good and bad choices for the frequency index set $\mathcal{A}\subseteq\mathds{Z}^{d}$ and the generating vector $\bm{z}\in\mathds{Z}^{d}$. The following theorem gives an error bound for the classical lattice algorithm and follows [CKNS20, Section 2.4].
@@
Let $H$ be the function space defined in (2.3), $\bm{X}=\{\tfrac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ a lattice, and $\mathcal{A}=\{\bm{h}\in\mathds{Z}^{d}:r(\bm{h})\leq M\}$ for some radius $M>0$. Then for the classical lattice algorithm defined in (2.1), we have
$e^{\mathrm{wor-app}}(A_{\mathcal{A}}^{\bm{X}})^{2}<\frac{1}{M}+M\mathcal{S}_{n}(\bm{z})\,,$ (2.4)
@@
$\mathcal{S}_{n}(\bm{z})=\sum_{\bm{h}\in\mathds{Z}^{d}}\frac{1}{r(\bm{h})}\sum_{\begin{subarray}{c}\bm{\ell}\in\mathds{Z}^{d}\setminus\{\bm{0}\}\\ \langle\bm{\ell},\bm{z}\rangle\equiv_{n}0\end{subarray}}\frac{1}{r(\bm{h}+\bm{\ell})}\,.$ (2.5)
In particular, with the optimal choice $M=1/\sqrt{\mathcal{S}_{n}(\bm{z})}$, which minimizes (2.4), we have
$e^{\mathrm{wor-app}}(A_{\mathcal{A}}^{\bm{X}})^{2}<\frac{2}{M}=2\sqrt{\mathcal{S}_{n}(\bm{z})}\,.$ (2.6)
@@
For weighted Korobov spaces (see Section 5 below) and a given lattice size $n\in\mathds{N}$, component-by-component (CBC) constructions for generating vectors $\bm{z}$ are known, controlling $\mathcal{S}_{n}(\bm{z})$ as in Theorem 5.1 below, which is [KMN24, Theorem 3.3]. Fast CBC implementations are given in [CKNS20, GS25].
Next we comment on the kernel method, which is often used in practice and benefits from the lattice structure, cf. [KKKNS21, Section 5]. For kernel methods we need to assume that $H$ is a reproducing kernel Hilbert space, i.e., function evaluations are continuous, which is a common and natural assumption. By the Moore–Aronszajn theorem $H$ has a corresponding reproducing kernel $K\colon\mathds{T}^{d}\times\mathds{T}^{d}\to\mathds{C}$, which is symmetric and positive definite. The kernel approximation of $f$ based on the lattice $\bm{X}$ is defined by
$A_{\text{ker}}^{\bm{X}}f\coloneqq\sum_{k=0}^{n-1}a_{k}\,K(\cdot,\tfrac{k}{n}\bm{z}\bmod{}1)\quad\text{with}\quad\bm{a}=[a_{k}]_{k\in\{0,\dots,n-1\}}=\bm{K}^{-1}\bm{f}\,,$ (2.7)
@@
$\bm{K}=\Big{[}K(\tfrac{k}{n}\bm{z}\bmod{}1,\tfrac{k^{\prime}}{n}\bm{z}\bmod{}1)\Big{]}_{k,k^{\prime}\in\{0,\dots,n-1\}}=\Big{[}K(\tfrac{k-k^{\prime}}{n}\bm{z}\bmod{}1,\bm{0})\Big{]}_{k,k^{\prime}\in\{0,\dots,n-1\}}\,.$
Thus, it is possible to diagonalize it by a discrete Fourier transform (DFT) matrix $\bm{F}=[\exp(2\pi\mathrm{i}\,k\,k^{\prime}/n)]_{k,k^{\prime}\in\{0,\dots,n-1\}}$
@@
In terms of the error it is known that for any given set of points, the kernel approximation (2.7) is optimal among all (linear or nonlinear) algorithms using function values at the same point set in the worst-case setting with respect to any error norm, cf. [KKKNS21]. In particular, we have that the worst-case error of the kernel approximation is bounded by the worst-case error of the classical lattice algorithm (2.1)
@@
Thus, the bound from Theorem 2.1 applies to the kernel method as well.
3 Projection property of approximation algorithms
@@
Clearly we have linearity for $A_{\mathcal{A}}^{\bm{X}}$ but the idempotence is not always present. For instance, if there are two aliasing frequencies, i.e., $\bm{h},\bm{h}^{\prime}\in\mathcal{A}$ with $\langle\bm{h},\bm{z}\rangle\equiv_{n}\langle\bm{h}^{\prime},\bm{z}\rangle$, we can take $f\coloneqq\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)+\exp(2\pi\mathrm{i}\langle\bm{h}^{\prime},\cdot\rangle)$. We then have $f\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{A}}$ yet by (2.2)
@@
A lattice $\bm{X}=\{\tfrac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ is said to have the reconstructing property on $\mathcal{B}\subseteq\mathds{Z}^{d}$ if
$\langle\bm{h},\bm{z}\rangle\not\equiv_{n}\langle\bm{h}^{\prime},\bm{z}\rangle\quad\text{for all}\quad\bm{h},\bm{h}^{\prime}\in\mathcal{B}\,.$ (3.1)
This is also sufficient for the idempotence as the following theorem shows. For ease of readability, we will write $\mathcal{B}\subseteq\mathds{Z}^{d}$ when we have the reconstructing property and $\mathcal{A}\subseteq\mathds{Z}^{d}$ for a generic frequency index set.
@@
By definition we have that $A_{\mathcal{B}}^{\bm{X}}$ is linear and maps onto $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$. Combining the aliasing formula (2.2) with the reconstructing property of $\bm{X}$ on $\mathcal{B}$ as in (3.1), we obtain $\widehat{g}_{\bm{h}}=\widehat{f}_{\bm{h}}$ for all $\bm{h}\in\mathcal{B}$ and confirm the idempotence. ∎
@@
Let $\bm{X}=\{\frac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ be a lattice. Further, let $H$ be the function space from (2.3) and $A^{\bm{X}}$ a (linear or non-linear) approximation algorithm using samples of $\bm{X}$. Then $\bm{X}$ is a lattice with reconstructing property (3.1) for the frequency index set
$\Big{\{}\bm{h}\in\mathds{Z}^{d}:r(\bm{h})<(e^{\mathrm{wor-app}}(A^{\bm{X}}))^{-2}\Big{\}}\,.$ (3.2)
In particular, the above holds for the best possible worst-case error achievable by any (linear or non-linear) algorithm using points from the lattice $\bm{X}$.
@@
$\Big{\{}\bm{h}\in\mathds{Z}^{d}:r(\bm{h})\leq\frac{1}{2\sqrt{\mathcal{S}_{n}(\bm{z})}}\Big{\}}\,.$ (3.3)
@@
Using the classical lattice algorithm (2.1) in Theorem 3.2, we obtain the reconstructing property on
@@
Now we plug in the bound (2.6), which only makes the frequency index set smaller, not losing the reconstructing property. We have $\leq$ insead of $<$ in (3.3) since equality is not attained in (2.1). ∎
Note, the frequency index set (3.3) has half of the radius of the frequency index set used in the classical lattice algorithm in Theorem 2.1, where we have $M=1/\sqrt{\mathcal{S}_{n}(\bm{z})}$. With the latter choice we do not have the guarantee for the reconstructing property. This difference could be an artifact of the proof, however.
4 Error bounds for subsampled rank-1 lattices

### v3

@@ Abstract.
2020 Mathematics Subject Classification:
41A25, 94A20
1. Introduction
This paper mainly deals with the question of approximating functions using a subsample of a rank-1 lattice. Rank-1 lattices have their origin as a Quasi-Monte Carlo (QMC) method and have been thoroughly studied in various settings ever since, see e.g., [Sloan94, kaemmererdiss, KaPoVo13, PlPoStTa18, CKNS20, DKP22]. The lattice point sets are given by $\bm{X}=\{\tfrac{k}{n}\bm{z}\bmod{}1\}_{k\in\{0,\dots,n-1\}}$, where $\bm{z}\in\mathds{Z}^{d}$ is known as the generating vector and it controls the quality of the point set, and $\bmod\,1$ denotes a component-wise fractional part of a vector. The study of partial or subsampled rank-1 lattices has only begun recently, cf. [BKPU22]. Such point sets have the form
@@ 1. Introduction
An example is depicted in Figure˜1. The structure of lattice point sets allows for the application of fast Fourier methods when approximating functions whilst the subsampling improves the approximation properties, cf. [BSU23]. We investigate the worst-case setting, where the point set and algorithm are fixed and are used to approximate every function from a given function class $H$. The worst-case error for a general algorithm $A$ measured in $L_{2}$-norm is defined by
@@
$\bm{X}$$\bm{X}_{J}$Refer to caption
Figure 1. Left: rank-1 lattice with generating vector $\bm{z}=(1,21)$ and lattice size $n=55$.
Right: subsampled rank-1 lattice with subsampling size $|J|=\lceil\sqrt{n}\,\log\sqrt{n}\,\rceil=15$.
@@
    (1.1) $e^{\mathrm{wor-app}}(A_{\mathcal{A}}^{\bm{X}})\lesssim n^{-\alpha/2+\varepsilon}\,$,
    where $\varepsilon>0$, $A_{\mathcal{A}}^{\bm{X}}$ is the “classical lattice algorithm” (see (2.1) below), and $A_{n}\lesssim B_{n}$ means there exists a constant $C>0$ independent of $n$ such that $A_{n}\leq CB_{n}$ for all $n\in\mathds{N}$. When $\alpha$ is an integer it is the number of available square-integrable derivatives in each coordinate.
@@
In this paper we connect both approaches by finding a frequency index set for which a lattice is reconstructing in Theorem˜3.3 below. These frequencies solely depend on the worst-case error achievable using any algorithm and the lattice in question. In that way, lattices constructed according to the first approach can be investigated with respect to their reconstructing property, and tools from the second approach become available.
@@
(1.2) $e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})\lesssim|J|^{-\alpha}(\log|J|)^{(d-1)\alpha+\frac{1}{2}}\quad\text{with}\quad n\sim|J|^{\frac{2}{1-1/(2\alpha)}}\,$,
where $S_{\mathcal{B}}^{\bm{X}_{J}}$ is the “least squares approximation” to be defined precisely in (4.1) below, where $J$ labels the subsampled point set $\bm{X}_{J}$ and $A\sim B$ means $A\lesssim B$ and $B\lesssim A$ holds simultaneously.
The method in [BKPU22] achieves the optimal polynomial order $|J|^{-\alpha+\varepsilon}$ up to $\varepsilon>0$ in terms of sampling complexity. However, the computational complexity is determined by $n$. Since an algorithm using a subsampled lattice is inherently still using the lattice of size $n$, the lower bound from [BKUV17] still applies, and achieving the rate $|J|^{-\alpha+\varepsilon}$ implies $n^{-\alpha/2}\lesssim|J|^{-\alpha+\varepsilon}$, yielding the lower bound
@@
With respect to the size of $|J|$, $n$ given in (1.2) is larger still, being off by a polynomial order depending on the smoothness $\alpha$.
In this paper, in Theorem˜5.3 and Corollary˜5.4 below, we show that we can go arbitrarily close to the squared number of points independently of the smoothness, i.e., we have
$e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})\lesssim|J|^{-\alpha+\varepsilon}\quad\text{with}\quad|J|^{2\sqrt{1-\varepsilon/\alpha}}\lesssim n\lesssim|J|^{2/\sqrt{1-\varepsilon/\alpha}}$
@@
This paper is organized as follows. We start by introducing the classical lattice algorithm and the kernel method in Section˜2. Then we investigate the projection property of approximation algorithms in Section˜3, where we also prove a connection between the worst-case error and the reconstructing property of lattices in Theorem˜3.3, which appears to be a new result. In Section˜4 we show general error bounds for the randomly subsampled lattice, which we apply to the Korobov space setting in Section˜5. After some comments on different implementations in Section˜6, we present some numerical experiments in Section˜7 and give concluding remarks in Section˜8.
@@
In this paper $\mathds{T}=\mathds{R}/\mathds{N}$ is the $1$-periodic torus; $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product; $\bm{x}\bmod{}1$ is understood as the component-wise fractional part of a vector $\bm{x}\in\mathds{R}^{d}$; we write $a\equiv_{n}b$ if there exists $k\in\mathds{N}$ such that $a=b+kn$ and $a\not\equiv_{n}b$ if not; $\|\cdot\|_{2}$ denotes the spectral norm; vectors are understood as column vectors.
2. Classical lattice algorithm versus kernel method
@@
(2.1) $A_{\mathcal{A}}^{\bm{X}}f\coloneqq\sum_{\bm{h}\in\mathcal{A}}\Big(\underbrace{\frac{1}{n}\sum_{k=0}^{n-1}f(\tfrac{k}{n}\bm{z}\bmod{}1)\exp(-2\pi\mathrm{i}\tfrac{k}{n}\langle\bm{h},\bm{z}\rangle)}_{\eqqcolon\,\widehat{g}_{\bm{h}}}\Big)\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\,$,
see e.g., [KSW06, KaPoVo13, kaemmererdiss, PlPoStTa18, CKNS20, DKP22]. To get a feeling for this method, we have a look at the Fourier coefficients of the approximation $A_{\mathcal{A}}^{\bm{X}}f\eqqcolon g=\sum_{\bm{h}\in\mathcal{A}}\widehat{g}_{\bm{h}}\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)$ in terms of the Fourier coefficients of $f$. By the character property
$\frac{1}{n}\sum_{k=0}^{n-1}\exp(2\pi\mathrm{i}m\tfrac{k}{n})=\delta_{m\equiv_{n}0}\coloneqq\begin{cases}1&\text{if }m\equiv_{n}0,\\ 0&\text{otherwise,}\end{cases}$
we have for $\bm{h}\in\mathcal{A}$,
@@
(2.2) $=\sum_{\begin{subarray}{c}\bm{\ell}\in\mathds{Z}^{d}\\ \langle\bm{\ell},\bm{z}\rangle\equiv_{n}\langle\bm{h},\bm{z}\rangle\end{subarray}}\!\!\!\!\!\!\widehat{f}_{\bm{\ell}}\,$.
@@
(2.3) $f\in H=\Big\{f\colon\mathds{T}^{d}\to\mathds{C}:\|f\|_{H}^{2}=\sum_{\bm{h}\in\mathds{Z}^{d}}r(\bm{h})\,|\widehat{f}_{\bm{h}}|^{2}<\infty\Big\}\,$,
where $r(\bm{h})\to\infty$ for $\|\bm{h}\|_{2}\to\infty$. Specific choices of the function $r(\cdot)$ lead to e.g., the Korobov spaces in Section˜5. Obviously there are good and bad choices for the frequency index set $\mathcal{A}\subseteq\mathds{Z}^{d}$ and the generating vector $\bm{z}\in\mathds{Z}^{d}$. The following theorem gives an error bound for the classical lattice algorithm and follows [CKNS20, Section 2.4].
@@
Let $H$ be the function space defined in (2.3). Consider a lattice $\bm{X}=\{\tfrac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ and a frequency index set $\mathcal{A}=\{\bm{h}\in\mathds{Z}^{d}:r(\bm{h})\leq M\}$ for some radius $M>0$. Then for the classical lattice algorithm defined in (2.1), we have
(2.4) $e^{\mathrm{wor-app}}(A_{\mathcal{A}}^{\bm{X}})^{2}<\frac{1}{M}+M\mathcal{S}_{n}(\bm{z})\,$,
@@
(2.5) $\mathcal{S}_{n}(\bm{z})=\sum_{\bm{h}\in\mathds{Z}^{d}}\frac{1}{r(\bm{h})}\sum_{\begin{subarray}{c}\bm{\ell}\in\mathds{Z}^{d}\setminus\{\bm{0}\}\\ \langle\bm{\ell},\bm{z}\rangle\equiv_{n}0\end{subarray}}\frac{1}{r(\bm{h}+\bm{\ell})}\,.$
In particular, with the optimal choice $M=1/\sqrt{\mathcal{S}_{n}(\bm{z})}$, which minimizes the right-hand side of (2.4), we have
(2.6) $e^{\mathrm{wor-app}}(A_{\mathcal{A}}^{\bm{X}})^{2}<\frac{2}{M}=2\sqrt{\mathcal{S}_{n}(\bm{z})}\,.$
@@
For weighted Korobov spaces (see Section˜5 below) and a given lattice size $n\in\mathds{N}$, component-by-component (CBC) constructions for generating vectors $\bm{z}$ are known, controlling $\mathcal{S}_{n}(\bm{z})$ as in Theorem˜5.1 below, which is [KMN24, Theorem 3.3]. Fast CBC implementations are given in [CKNS20, GS25].
Next we comment on the kernel method, which is often used in practice and benefits from the lattice structure, cf. [KKKNS21, Section 5]. For kernel methods we need to assume that $H$ is a reproducing kernel Hilbert space, i.e., function evaluations are continuous, which is a common and natural assumption. By the Moore–Aronszajn theorem, $H$ has a corresponding reproducing kernel $K\colon\mathds{T}^{d}\times\mathds{T}^{d}\to\mathds{C}$, which is symmetric and positive definite. The kernel approximation of $f$ based on the lattice $\bm{X}$ is defined by
(2.7) $A_{\text{ker}}^{\bm{X}}f\coloneqq\sum_{k=0}^{n-1}a_{k}\,K(\cdot,\tfrac{k}{n}\bm{z}\bmod{}1)\quad\text{with}\quad\bm{a}=[a_{k}]_{k\in\{0,\dots,n-1\}}=\bm{K}^{-1}\bm{f}\,$,
@@
$\bm{K}$ $=\Big[K(\tfrac{k}{n}\bm{z}\bmod{}1,\tfrac{k^{\prime}}{n}\bm{z}\bmod{}1)\Big]_{k,k^{\prime}\in\{0,\dots,n-1\}}$
$=\Big[K(\tfrac{k-k^{\prime}}{n}\bm{z}\bmod{}1,\bm{0})\Big]_{k,k^{\prime}\in\{0,\dots,n-1\}}\,.$
Thus, it is possible to diagonalize it by a discrete Fourier transform (DFT) matrix $\bm{F}=[\exp(2\pi\mathrm{i}\,k\,k^{\prime}/n)]_{k,k^{\prime}\in\{0,\dots,n-1\}}$ as
@@
In terms of the error it is known that for any given set of points, the kernel approximation (2.7) is optimal among all (linear or nonlinear) algorithms using function values at the same point set in the worst-case setting with respect to any error norm, cf. [KKKNS21]. In particular, we have that the worst-case error of the kernel approximation is bounded by the worst-case error of the classical lattice algorithm (2.1), i.e.,
@@
Thus, the bound from Theorem˜2.1 applies to the kernel method as well.
3. Projection property of approximation algorithms
@@
Clearly we have linearity for $A_{\mathcal{A}}^{\bm{X}}$ but the idempotence is not always present. For instance, if there are two aliasing frequencies, i.e., $\bm{h},\bm{h}^{\prime}\in\mathcal{A}$ with $\langle\bm{h},\bm{z}\rangle\equiv_{n}\langle\bm{h}^{\prime},\bm{z}\rangle$, we can take $f\coloneqq\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)+\exp(2\pi\mathrm{i}\langle\bm{h}^{\prime},\cdot\rangle)$. We then have $f\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{A}}$ and yet by (2.2) we have
@@
Definition 3.1.
A lattice $\bm{X}=\{\tfrac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ is said to have the reconstructing property on a frequency index set $\mathcal{B}\subseteq\mathds{Z}^{d}$ if
(3.1) $\langle\bm{h},\bm{z}\rangle\not\equiv_{n}\langle\bm{h}^{\prime},\bm{z}\rangle\quad\text{for all}\quad\bm{h},\bm{h}^{\prime}\in\mathcal{B}\,.$
The reconstructing property is also sufficient for the idempotence as the following theorem shows. For ease of readability, we will write $\mathcal{B}\subseteq\mathds{Z}^{d}$ when we have the reconstructing property and $\mathcal{A}\subseteq\mathds{Z}^{d}$ for a generic frequency index set.
@@
By definition we have that $A_{\mathcal{B}}^{\bm{X}}$ is linear and maps onto the linear space $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$. Combining the aliasing formula (2.2) with the reconstructing property of $\bm{X}$ on $\mathcal{B}$ as in (3.1), we obtain $\widehat{g}_{\bm{h}}=\widehat{f}_{\bm{h}}$ for all $\bm{h}\in\mathcal{B}$ and confirm the idempotence. ∎
@@
Let $\bm{X}=\{\frac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ be a lattice. Further, let $H$ be the function space from (2.3) and $A^{\bm{X}}$ a (linear or non-linear) approximation algorithm using only samples from $\bm{X}$. Then $\bm{X}$ is a lattice with reconstructing property (3.1) for the frequency index set
(3.2) $\Big\{\bm{h}\in\mathds{Z}^{d}:r(\bm{h})<(e^{\mathrm{wor-app}}(A^{\bm{X}}))^{-2}\Big\}\,.$
In particular, the reconstructing property holds for the index set (3.2) defined with the best possible worst-case error achievable by any (linear or non-linear) algorithm using points from the lattice $\bm{X}$.
@@
(3.3) $\Big\{\bm{h}\in\mathds{Z}^{d}:r(\bm{h})\leq\frac{1}{2\sqrt{\mathcal{S}_{n}(\bm{z})}}\Big\}\,.$
@@
Using the classical lattice algorithm (2.1) in Theorem˜3.3, we obtain the reconstructing property on
@@
Now we plug in the bound (2.6), which only makes the frequency index set smaller. By definition, $\bm{X}$ retains the reconstructing property on the smaller set. We have $\leq$ instead of << in (3.3) since equality is not attained in (2.1). ∎
Note, the frequency index set (3.3) has half of the radius of the frequency index set used in the classical lattice algorithm in Theorem˜2.1, where we have $M=1/\sqrt{\mathcal{S}_{n}(\bm{z})}$. With the latter choice we do not have the guarantee for the reconstructing property. This difference could be an artifact of the proof, however.
4. Error bounds for subsampled rank-1 lattices